# CFSP: An Efficient Structured Pruning Framework for LLMs with Coarse-to-Fine Activation Information

### v1

@@ Abstract
Yuxin Wang [^1] footnotemark: 1, Minghua Ma [^1] footnotemark: 1, Zekun Wang1thanks: Equal Contribution.thanks: Corresponding Author., Jingchang Chen1, Huiming Fan1 Liping Shan2, Qing Yang2, Dongliang Xu2, Ming Liu122footnotemark: 2, Bing Qin1 1Harbin Institute of Technology, Harbin, China 2Du Xiaoman Science Technology Co., Ltd, Beijing, China {yuxinwang,mhma,zkwang,jcchen,hmfan,mliu,qinb}@ir.hit.edu.cn
@@ 1 Introduction
Although scaling up Large Language Models (LLMs) brings remarkable performance (Brown et al., 2020; OpenAI, 2023; Gemini Team et al., 2023; Meta, 2024; DeepSeek-AI et al., 2024; Yang et al., 2024a), increasing parameters brings more computations and memory consumption, posing a significant challenge of deploying in practical applications. To address this, various model compression methods for LLMs are proposed (Dettmers et al., 2022; Frantar et al., 2022; Lin et al., 2023; Muralidharan et al., 2024). Existing LLM pruning work (Frantar and Alistarh, 2023; Sun et al., 2023; Xu et al., 2024a; Zhang et al., 2024b) focuses mainly on unstructured or semi-structured sparsity. However, these paradigms require specific hardware to achieve practical acceleration.
@@
To this end, we propose CFSP (shown in Figure 1), an efficient structural pruning framework for LLM that takes advantage of coarse to fine-grained activation information to guide pruning. Specifically, we employ activation as the importance criterion, which is calculated for blocks (coarse-grained) and the weights within each block (fine-grained) in a single forward pass. For each block, we measure its saliency of transformations based on the angular distance of the input and output representations. Then, we utilize this criterion as coarse-grained importance to assign the sparsity budget. Finally, for weights within each block, we take the product of their relative activations and weights as a fine-grained criterion to remove redundant parts. Since existing work typically performs recovery fine-tuning with LoRA to further improve performance, we propose a more efficient recovery method that leverages coarse-grained importance to adaptively allocate additional trainable parameters: the pruned models can achieve comparable performance while utilizing less recovery data.
Our contributions can be summarized as:
@@
  • We introduce an efficient recovery fine-tuning method that adaptively assigns additional trainable parameters based on coarse-grained importance score.
@@ 2.2 CFSP Framework
CFSP takes activations as an importance criterion to identify redundant parts of LLMs for the following reasons: (1) Activations can be obtained with a single forward pass, resulting in significantly lower overhead compared to other metrics. (2) As pointed out in previous studies (Sun et al., 2023; Lin et al., 2023), parameter weights corresponding to larger activation magnitudes are more salient since they process more important features.
@@
The Equation 5 may assign irregularly shaped weight matrices that do not satisfy the multiples of 64 or 128, thus destroying the parallelism of the tensor cores on the GPU. To this end, we introduce a simple adjustment during pruning to adjust the final dimensions of the pruned blocks to multiples of 128. For the $l$-th block $\mathbf{B}^{l}$, the final dimension $dim_{f}^{l}$ are computed as:
$dim_{f}^{l}$ $=(\left\lfloor\frac{dim_{o}\times\operatorname{Sparsity}(\mathbf{B}^{i})+64}{128}\right\rfloor)\times 128$ (6)
@@
After assigning the proper sparsity to each block, we then identify the importance of pruning units (intermediate dimensions) within the block. Figure 3 shows the structural dependencies of three matrices used in FFN ($W_{u}$, $W_{g}$ and $W_{d}$): removing the intermediate dimensions is equivalent to subtracting the corresponding columns of $W_{u},W_{g}$ and the corresponding rows of $W_{d}$. The weight matrix represents the connections between neurons, where each row or column of weights influences the same neuron, implying that the fine-grained importance of weights is also related to their respective row or column. For the $i$-th intermediate dimension, we utilize the activation of $\mathbf{X}_{d}$ of $W_{d}$ and all weight matrices to calculate fine-grained importance score $S_{l}^{i}$:
@@
$F_{l}^{i}$ $=\sum_{j}(\frac{W^{ij}_{d}\cdot\left\|{\mathbf{X}}^{i}_{d}\right\|}{W^{*j}_{d}\cdot\left\|{\mathbf{X}}^{*}_{d}\right\|}+\frac{W^{ij}_{u}}{W^{i*}_{u}}+\frac{W^{ij}_{g}}{W^{i*}_{g}})$ (8)
@@ 2.3 Importance-guided Recovery Fine-tuning
where $\bar{r}$ is the average rank allocated budget. In our experiments, we find that our recovery method is more efficient, achieving comparable performance while requiring less training data.
@@ 3.1 Experimental Setup
Following previous work (Sun et al., 2023; An et al., 2024), we evaluate the zero-shot performance of models across 5 well-known tasks: WinoGrande (Sakaguchi et al., 2020), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018). Since Jaiswal et al. (2023a) shows that LLM pruning methods tend to have significantly degraded performance on knowledge-intensive tasks, we also include two challenging QA tasks for zero-shot evaluation: MMLU (Hendrycks et al., 2021) and FreebaseQA (Jiang et al., 2019), which focus on factual knowledge. For language modeling performance, we evaluate models on WikiText2 (Merity et al., 2017). Following previous work, we use the LM-Evaluation-Harness (Gao et al., 2023) and LLM-Kick (Jaiswal et al., 2023a) with default hyperparameters for the corresponding tasks. More details of the evaluation are shown in Appendix A.1
@@
For the pruning stage, the calibration data is randomly selected from the WikiText2 (Merity et al., 2017) training set. Unless otherwise stated, the calibration set consists of 128 samples and each has approximately 1024 tokens following Sun et al. (2023). In the recovery fine-tuning stage, the pruned models are trained on the 0.1B tokens from FineWeb-Edu (Lozhkov et al., 2024) dataset with the next-token prediction loss. We set the average rank budget of IG-LoRA at 8 following Ma et al. (2023). More details and ablation of the implementation are shown in Appendix A.2 and Appendix B.2, respectively.
@@
We compare the single-shot pruning performance [^2] Without recovery fine-tuning.of CFSP against the following baselines: Magnitude-SP measures the importance criterion based on the magnitude of weights (Han et al., 2016; Jaiswal et al., 2023b). This baseline employs uniform sparsity across blocks. Wanda-SP is extended by the unstructured pruning method Wanda (Sun et al., 2023), which modifies the target pruning units to structured weights. We globally sort the pruning units across all blocks to identify redundant components, as this strategy tends to achieve better performance compared to adopting a local manner for individual blocks. FLAP (An et al., 2024) uses the stability of activations as an importance criterion, also applying a global sorting strategy. Notably, for a fair comparison, all baselines are implemented to prune intermediate dimensions of the FFN, which are the same as CFSP. Details are shown in Appendix A.3.
@@ 3.2 Main Results
We present a performance comparison of the LLaMA3 family in Tables 1 and 2, as well as LLaMA2-13B in Table  3. In the single-shot pruning setting, CFSP consistently demonstrates superior average performance compared to baselines across various models at both 20% and 50% sparsity. Remarkably, CFSP achieves a promising accuracy of 32.39 on MMLU with 50% sparsity on LLaMA3-8B, while other baselines regress to chance-level accuracy (~25.0). This result underscores the potential of CFSP to perform well on more challenging tasks without retraining, even at high sparsity. Furthermore, CFSP is more favorable for larger models. At the 20% and 50% sparsity on LLaMA3-70B, CFSP maintains 97.9% and 82.2% of the original performance on average, respectively. We further experiment CFSP with recovery fine-tuning at 50% sparsity for each model. For comparison, we choose Wanda-SP, as it has the second-best average performance in single-shot pruning. We fine-tune pruned models with our proposed IG-LoRA on 0.1B tokens from the FineWeb-Edu dataset. We find that after recovery training, both pruning models are improved, especially on complex knowledge-sensitive tasks. CFSP still outperforms Wanda-SP in general, indicating the effectiveness of our proposed pruning and recovery approach.
@@ 3.4 Ablation Study
We explore the effects of each component incorporated in the importance criterion of the proposed CFSP. Table 6 shows the ablation results under 20% sparsity of LLaMA3-8B. We first investigate the coarse-grained importance of blocks by comparing variants including: (a) uniform sparsity for each block, (b) Euclidean distance, or (c) cosine similarity as the coarse-grained importance criterion to allocation sparsity budget across blocks. As illustrated in Table 6, applying uniform sparsity or using Euclidean distance results in a notable performance decrease, particularly for zero-shot tasks. The angular distance (Eq. 2) used in CFSP achieves the best performance across tasks. For fine-grained importance ablation, as shown in the Table 6, the criterion outlined in Eq. 8 also demonstrates superior performance compared to the criterion utilized in Wanda.
@@ 4 Related Work
The enormous computations of LLMs has prompted efforts in improving their efficiency, including quantization (Dettmers et al., 2022; Yao et al., 2022; Lin et al., 2023; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023a, b; Shao et al., 2023; Xu et al., 2024b), distillation (Wang et al., 2022; Jiang et al., 2023; Wang et al., 2023; Hsieh et al., 2023; Agarwal et al., 2023; Gu et al., 2023) and KV cache compression (Sheng et al., 2023; Ge et al., 2023; Zhang et al., 2023b; Liu et al., 2024; Hooper et al., 2024; DeepSeek-AI et al., 2024). Pruning is another crucial method by eliminating redundant parameters. Most of the previous pruning work follows the unstructured pruning, which removes individual parameters according to their importance (Frantar and Alistarh, 2023; Sun et al., 2023; Dettmers et al., 2023b; Xu et al., 2024a; Zhang et al., 2024b). However, this paradigm requires specialized hardware support to speed up. In contrast, structured pruning eliminates the structural group of weights, facilitating a more convenient deployment on general hardware (Wang et al., 2020; Xia et al., 2022). Some work proposes to remove redundant layers in LLMs (Men et al., 2024; Yang et al., 2024b; Gromov et al., 2024b), while dropping entire layers leads to a significant performance drop. For pruning on more fine-grained units, some work formulates pruning as a constrained optimization problem by introducing learnable masks to search (Xia et al., 2023; Dery et al., 2024; Muralidharan et al., 2024; Gunter et al., 2024). Zhang et al. (2024a) performs iteratively to prune the coupled weights until the desired sparsity is achieved. Ma et al. (2023) and Zhang et al. (2023a) use gradient information to guide pruning. These methods incur substantial pruning overhead, particularly in the case of large-scale models. An et al. (2024) eliminates channels based on their activation fluctuations using only forward passes. In this work, we also aim to achieve efficient structured pruning using only forward passes.
@@ A.1 Details of Evaluation Benchmarks
In this work, we consider the following tasks for evaluating zero-shot performance, along with their respective evaluation metrics: WinoGrande (Sakaguchi et al., 2020) with the accuracy, PIQA (Bisk et al., 2020) with the normalized accuracy, OBQA (Mihaylov et al., 2018) with the normalized accuracy, HellaSwag (Zellers et al., 2019) with the normalized accuracy, ARC-easy/challenge (Clark et al., 2018) with the normalized accuracy, MMLU (Hendrycks et al., 2021) with the accuracy, FreebaseQA (Jiang et al., 2019) with the exact-match score. The first 6 tasks are general common sense reasoning tasks, while the others are knowledge-intensive. We evaluate WinoGrande, PIQA, HellaSwag, BoolQ, ARC-e/c, and MMLU by LM-Evaluation-Harness (Gao et al., 2023) in multiple choice form: we compute the loglikelihood for each choice and report the accuracy for the highest choice. For FreebaseQA, the evaluation is run with LLM-Kick (Jaiswal et al., 2023a).
@@ A.3 Details of Baselines
measures the importance criterion based on the magnitude of weights (Han et al., 2016; Jaiswal et al., 2023b). This baseline employs with uniform sparsity across blocks.
@@
is extended by the unstructured pruning method Wanda (Sun et al., 2023), which modifies the target pruning units to structured weights. This baseline uses the product of weights and activations as an importance criterion. We globally sort the pruning units across all blocks to identify redundant components, as this strategy tends to achieve better performance compared to adopting a local manner for individual blocks.
@@
0% LLaMA2-7B 69.06 79.11 44.20 76.02 74.62 46.33 41.25 68.39 62.37
20% w/o recovery
Magnitude-SP 48.70 52.12 24.40 28.77 30.18 24.32 25.84 0.55 29.36
Wanda-SP 66.93 76.50 41.80 70.82 64.48 38.57 32.19 44.44 54.46
FLAP 65.51 75.84 40.00 69.64 60.82 37.29 30.86 28.65 51.07
CFSP (ours) 67.25 76.88 40.60 72.05 68.77 41.47 36.33 38.99 55.29
[…]
Magnitude-SP 50.20 48.20 27.00 26.32 26.52 29.10 26.84 0.53 29.34
Wanda-SP 61.56 66.49 34.80 52.23 45.37 28.07 25.45 4.15 39.76
FLAP 57.62 66.00 32.80 48.09 40.07 27.39 23.04 0.90 36.99
CFSP (ours) 61.64 67.36 35.20 53.96 48.61 30.20 23.07 4.35 40.55
w/ recovery
Wanda-SP 63.85 71.11 37.80 61.40 57.08 35.04 26.32 20.90 46.69
CFSP (ours) 65.11 70.73 37.00 61.96 58.88 36.26 29.46 20.05 47.43

### v2

@@ Abstract
Yuxin Wang [^1] footnotemark: 1, Minghua Ma [^1] footnotemark: 1, Zekun Wang1thanks: Equal Contribution.thanks: Corresponding Author., Jingchang Chen1, Huiming Fan1, Liping Shan2, Qing Yang2, Dongliang Xu2, Ming Liu122footnotemark: 2, Bing Qin1 1Harbin Institute of Technology, Harbin, China 2Du Xiaoman Science Technology Co., Ltd, Beijing, China {yuxinwang,mhma,zkwang,jcchen,mliu,qinb}@ir.hit.edu.cn
@@ 1 Introduction
Although scaling up Large Language Models (LLMs) brings remarkable performance (Brown et al., 2020; OpenAI, 2023; Gemini Team et al., 2023; Meta, 2024; DeepSeek-AI et al., 2024; Yang et al., 2024a), increasing parameters brings more computations and memory consumption, posing a significant challenge of deploying in practical applications. To address this, various model compression methods for LLMs are proposed (Dettmers et al., 2022; Frantar et al., 2022; Lin et al., 2024; Muralidharan et al., 2024). Existing LLM pruning work (Frantar and Alistarh, 2023; Sun et al., 2024; Xu et al., 2024a; Zhang et al., 2024b) focuses mainly on unstructured or semi-structured sparsity. However, these paradigms require specific hardware to achieve practical acceleration.
@@
To this end, we propose CFSP (shown in Figure 1), an efficient structural pruning framework for LLM that takes advantage of coarse to fine-grained activation information to guide pruning. Specifically, we employ activation as the importance criterion, which is calculated for blocks (coarse-grained) and the weights within each block (fine-grained) in a single forward pass. For each block, we measure its saliency of transformations on the basis of the angular distance of the input and output representations. Then, we utilize this criterion as coarse-grained importance to assign the sparsity budget. Finally, for weights within each block, we take the product of their relative activations and weights as a fine-grained criterion to remove redundant parts. Since existing work typically performs recovery fine-tuning with LoRA to further improve performance, we propose a more efficient recovery method that leverages coarse-grained importance to adaptively allocate additional trainable parameters: the pruned models can achieve comparable performance while utilizing less recovery data.
Our contributions can be summarized as follows:
@@
  • We introduce an efficient recovery fine-tuning method that adaptively assigns additional trainable parameters based on the coarse-grained importance score.
@@ 2.2 CFSP Framework
CFSP takes activations as an importance criterion to identify redundant parts of LLMs for the following reasons: (1) Activations can be obtained with a single forward pass, resulting in significantly lower overhead compared to other metrics. (2) As pointed out in previous studies (Sun et al., 2024; Lin et al., 2024), parameter weights corresponding to larger activation magnitudes are more salient since they process more important features.
@@
Equation 5 can assign irregularly shaped weight matrices that do not satisfy the multiples of 64 or 128, thus destroying the parallelism of the tensor cores on the GPU. To this end, we introduce a simple adjustment during pruning to adjust the final dimensions of the pruned blocks to multiples of 128. For the $l$-th block $\mathbf{B}^{l}$, the final dimension $dim_{f}^{l}$ are computed as:
$dim_{f}^{l}$ $=\left(\left\lfloor\frac{dim_{o}\times\operatorname{Sparsity}(\mathbf{B}^{i})+64}{128}\right\rfloor\right)\times 128$ (6)
@@
After assigning the proper sparsity to each block, we then identify the importance of pruning units (intermediate dimensions) within the block. Figure 3 shows the structural dependencies of three matrices used in FFN ($W_{u}$, $W_{g}$ and $W_{d}$): removing the intermediate dimensions is equivalent to subtracting the corresponding columns of $W_{u},W_{g}$ and the corresponding rows of $W_{d}$. The weight matrix represents the connections between neurons, where each row or column of weights influences the same neuron, implying that the fine-grained importance of the weights is also related to their respective row or column. For the $i$-th intermediate dimension, we utilize the activation of $\mathbf{X}_{d}$ of $W_{d}$ and all weight matrices to calculate fine-grained importance score $S_{l}^{i}$:
@@
$F_{l}^{i}$ $=\sum_{j}\left(\frac{W^{ij}_{d}\cdot\left\|{\mathbf{X}}^{i}_{d}\right\|}{W^{*j}_{d}\cdot\left\|{\mathbf{X}}^{*}_{d}\right\|}+\frac{W^{ij}_{u}}{W^{i*}_{u}}+\frac{W^{ij}_{g}}{W^{i*}_{g}}\right)$ (8)
@@ 2.3 Importance-guided Recovery Fine-tuning
where $\bar{r}$ is the averaged rank allocated budget. In our experiments, we find that our recovery method is more efficient, achieving comparable performance while requiring less training data.
@@ 3.1 Experimental Setup
Following previous work (Sun et al., 2024; An et al., 2024), we evaluate the zero-shot performance of models across 5 well-known tasks: WinoGrande (Sakaguchi et al., 2020), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018). Since Jaiswal et al. (2024) shows that LLM pruning methods tend to have significantly degraded performance on knowledge-intensive tasks, we also include two challenging QA tasks for zero-shot evaluation: MMLU (Hendrycks et al., 2021) and FreebaseQA (Jiang et al., 2019), which focus on factual knowledge. For language modeling performance, we evaluate models on WikiText2 (Merity et al., 2017). Following previous work, we use the LM-Evaluation-Harness (Gao et al., 2023) and LLM-Kick (Jaiswal et al., 2024) with default hyperparameters for the corresponding tasks. More details of the evaluation are shown in Appendix A.1
@@
For the pruning stage, the calibration data are randomly selected from the WikiText2 (Merity et al., 2017) training set. Unless otherwise stated, the calibration set consists of 128 samples and each has approximately 1024 tokens following Sun et al. (2024). In the recovery fine-tuning stage, pruned models are trained on 0.1B tokens from the FineWeb-Edu (Lozhkov et al., 2024) dataset with the next-token prediction loss. We set the average rank budget of IG-LoRA at 8 following Ma et al. (2023). More details and ablation of the implementation are shown in Appendix A.2 and Appendix B.2, respectively.
@@
We compare the single-shot pruning performance [^2] Without recovery fine-tuning.of CFSP against the following baselines: Magnitude-SP measures the importance criterion based on the magnitude of weights (Han et al., 2016; Jaiswal et al., 2023). This baseline employs uniform sparsity across blocks. Wanda-SP is extended by the unstructured pruning method Wanda (Sun et al., 2024), which modifies the target pruning units to structured weights. We globally sort the pruning units across all blocks to identify redundant components, as this strategy tends to achieve better performance compared to adopting a local manner for individual blocks. FLAP (An et al., 2024) uses the stability of activations as an importance criterion, also applying a global sorting strategy. Notably, for a fair comparison, all baselines are implemented to prune the intermediate dimensions of the FFN, which are the same as CFSP. Details are shown in Appendix A.3.
@@ 3.2 Main Results
We present a performance comparison of the LLaMA3 family in Tables 1 and 2, as well as LLaMA2-13B in Table  3. In the single-shot pruning setting, CFSP consistently demonstrates superior average performance compared to baselines across various models at both 20% and 50% sparsity. Remarkably, CFSP achieves a promising accuracy of 32.39 on MMLU with 50% sparsity on LLaMA3-8B, while other baselines regress to chance-level accuracy (~25.0). This result underscores the potential of CFSP to perform well on more challenging tasks without retraining, even at high sparsity. Furthermore, CFSP is more favorable for larger models. At the 20% and 50% sparsity on LLaMA3-70B, CFSP maintains 97.9% and 82.2% of the original performance on average, respectively. We further evaluate CFSP with recovery fine-tuning at 50% sparsity for each model. For comparison, we choose Wanda-SP, as it has the second-best average performance in single-shot pruning. We fine-tune pruned models with our proposed IG-LoRA on 0.1B tokens from the FineWeb-Edu dataset. We find that after recovery training, both pruning models are improved, especially on complex knowledge-sensitive tasks. CFSP still outperforms Wanda-SP in general, indicating the effectiveness of our proposed pruning and recovery approach.
@@ 3.4 Ablation Study
We explore the effects of each component incorporated in the importance criterion of the proposed CFSP. Table 6 shows the ablation results under 20% sparsity of LLaMA3-8B. We first investigate the coarse-grained importance of blocks by comparing variants including: (a) uniform sparsity for each block, (b) Euclidean distance, or (c) cosine similarity as the coarse-grained importance criterion to allocation sparsity budget across blocks. As illustrated in Table 6, applying uniform sparsity or using Euclidean distance results in a notable performance decrease, particularly for zero-shot tasks. The angular distance (Eq. 2) used in CFSP achieves the best performance across tasks. For fine-grained importance ablation, as shown in Table 6, the criterion outlined in Eq. 8 also demonstrates superior performance compared to the criterion utilized in Wanda.
@@ 4 Related Work
The enormous computations of LLMs has prompted efforts in improving their efficiency, including quantization (Dettmers et al., 2022; Yao et al., 2022; Lin et al., 2024; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023a, b; Shao et al., 2023; Xu et al., 2024b), distillation (Wang et al., 2022; Jiang et al., 2023; Wang et al., 2023; Hsieh et al., 2023; Agarwal et al., 2023; Gu et al., 2023) and KV cache compression (Sheng et al., 2023; Ge et al., 2023; Zhang et al., 2023b; Liu et al., 2024; Hooper et al., 2024; DeepSeek-AI et al., 2024). Pruning is another crucial method by eliminating redundant parameters. Most of the previous pruning work follows the unstructured pruning, which removes individual parameters according to their importance (Frantar and Alistarh, 2023; Sun et al., 2024; Dettmers et al., 2023b; Xu et al., 2024a; Zhang et al., 2024b). However, this paradigm requires specialized hardware support to speed up. In contrast, structured pruning eliminates the structural group of weights, facilitating a more convenient deployment on general hardware (Wang et al., 2020; Xia et al., 2022). Some work proposes to remove redundant layers in LLMs (Men et al., 2024; Yang et al., 2024b; Gromov et al., 2024b), while dropping entire layers leads to a significant performance drop. For pruning on more fine-grained units, some work formulates pruning as a constrained optimization problem by introducing learnable masks to search (Xia et al., 2023; Dery et al., 2024; Muralidharan et al., 2024; Gunter et al., 2024). Zhang et al. (2024a) performs iteratively to prune the coupled weights until the desired sparsity is achieved. Ma et al. (2023) and Zhang et al. (2023a) use gradient information to guide pruning. These methods incur substantial pruning overhead, particularly in the case of large-scale models. An et al. (2024) eliminates channels based on their activation fluctuations using only forward passes. In this work, we also aim to achieve efficient structured pruning using only forward passes.
@@ 5 Conclusion
Acknowledgment
The research in this article is supported by the National Science Foundation of China (U22B2059, 62276083), the Human-Machine Integrated Consultation System for Cardiovascular Diseases (2023A003).
@@ A.1 Details of Evaluation Benchmarks
In this work, we consider the following tasks for evaluating zero-shot performance, along with their respective evaluation metrics: WinoGrande (Sakaguchi et al., 2020) with the accuracy, PIQA (Bisk et al., 2020) with the normalized accuracy, OBQA (Mihaylov et al., 2018) with the normalized accuracy, HellaSwag (Zellers et al., 2019) with the normalized accuracy, ARC-easy/challenge (Clark et al., 2018) with the normalized accuracy, MMLU (Hendrycks et al., 2021) with the accuracy, FreebaseQA (Jiang et al., 2019) with the exact-match score. The first 6 tasks are general common sense reasoning tasks, while the others are knowledge-intensive. We evaluate WinoGrande, PIQA, HellaSwag, BoolQ, ARC-e/c, and MMLU by LM-Evaluation-Harness (Gao et al., 2023) in multiple choice form: we compute the loglikelihood for each choice and report the accuracy for the highest choice. For FreebaseQA, the evaluation is run with LLM-Kick (Jaiswal et al., 2024).
@@ A.3 Details of Baselines
measures the importance criterion based on the magnitude of weights (Han et al., 2016; Jaiswal et al., 2023). This baseline employs with uniform sparsity across blocks.
@@
is extended by the unstructured pruning method Wanda (Sun et al., 2024), which modifies the target pruning units to structured weights. This baseline uses the product of weights and activations as an importance criterion. We globally sort the pruning units across all blocks to identify redundant components, as this strategy tends to achieve better performance compared to adopting a local manner for individual blocks.
@@
Sparsity Method Average PPL
0% Qwen2.5-7B 68.58 7.64
20% w/o recovery
Wanda-SP 62.70 8.82
FLAP 61.71 9.12
CFSP (ours) 63.02 9.03
Table 8: The averaged zero-shot performance and PPL on wikitext2 of pruned models on Qwen2.5-7B under 20% sparsity. Bold indicates the best results.
@@
0% LLaMA2-7B 69.06 79.11 44.20 76.02 74.62 46.33 41.25 68.39 62.37
20% w/o recovery
Magnitude-SP 48.70 52.12 24.40 28.77 30.18 24.32 25.84 0.55 29.36
Wanda-SP 66.93 76.50 41.80 70.82 64.48 38.57 32.19 44.44 54.46
FLAP 65.51 75.84 40.00 69.64 60.82 37.29 30.86 28.65 51.07
CFSP (ours) 67.25 76.88 40.60 72.05 68.77 41.47 36.33 38.99 55.29
[…]
Magnitude-SP 50.20 48.20 27.00 26.32 26.52 29.10 26.84 0.53 29.34
Wanda-SP 61.56 66.49 34.80 52.23 45.37 28.07 25.45 4.15 39.76
FLAP 57.62 66.00 32.80 48.09 40.07 27.39 23.04 0.90 36.99
CFSP (ours) 61.64 67.36 35.20 53.96 48.61 30.20 23.07 4.35 40.55
w/ recovery
Wanda-SP 63.85 71.11 37.80 61.40 57.08 35.04 26.32 20.90 46.69
CFSP (ours) 65.11 70.73 37.00 61.96 58.88 36.26 29.46 20.05 47.43